OpenStudy (anonymous):

find the solutions to the system. y = x^2 - 3x - 1 y=8x - 1 a. (0,-1) and (11,388) b. (0,01) and (11,87) c. (-1,0) and (87,11) d. no solutions

OpenStudy (bee_see):

replace x and y with the coordinates given and see which ones work

OpenStudy (jhannybean):

plug in \(y=8x-1\) into \(y\) and solve for \[8x-1=x^2-3x-1\]

OpenStudy (jhannybean):

So you move the function on the LHS to the right and set it = 0. \[0=x^2+x(-3-8)-1+1\]

OpenStudy (anonymous):

is it a ? @Jhannybean

OpenStudy (jhannybean):

Can you show me how you got a?